Hel CH 4 HW 0.60 points based costing system. In for the possible switch, Bunker has identfied two activity cost pools: materials handling and setup. Pertinent data follow 780 70
Answer #1
1.) Calculation of overhead cost assigned to leather case using traditional costing:
Total labor hours= 16000+9000= 25000
Overhead cost= 194500
Overhead cost per hour= overhead cost/ labor hours
                                                = 194500/25000=7.78
Overhead cost allocated to leather case= 9000*7.78= $70020
2) Calculation of overhead cost assigned to leather case using ABC:
(1)Calculation of activity rate
Activity cost pools Estimated overhead(a) Estimated cost driver(b) activity rates((a)/(b))
Machine moves 1,62,500 1300 125.00
Setups 32,000 100 320.00
Calculation of total overhead cost:
Activity cost pools (a) cost driver activity rates(b) cost(a*b)(in$)
Machine moves 780 125.00 97500.00
Setups 70 320.00 22400.00
Overhead cost allocated to leather case 119900.00
3) Overhead cost is underapplied in case of traditional costing as compared to ABC.